# Following the TRACE: A Structured Path to Empathetic Response Generation with Multi-Agent Models

### v1

@@ Abstract
Empathetic response generation is a crucial task for creating more human-like and supportive conversational agents. However, existing methods face a core trade-off between the analytical depth of specialized models and the generative fluency of Large Language Models (LLMs). To address this, we propose TRACE, Task-decomposed Reasoning for Affective Communication and Empathy, a novel framework that models empathy as a structured cognitive process by decomposing the task into a pipeline for analysis and synthesis. By building a comprehensive understanding before generation, TRACE unites deep analysis with expressive generation. Experimental results show that our framework significantly outperforms strong baselines in both automatic and LLM-based evaluations, confirming that our structured decomposition is a promising paradigm for creating more capable and interpretable empathetic agents. Our code is available at https://anonymous.4open.science/r/TRACE-18EF/README.md.
@@ 2.1 Modeling Empathetic Understanding
Early research in Empathetic Response Generation primarily focused on modeling the user’s inner state to build a foundation for empathetic interaction [6]. This pursuit branched into two main research streams, affective perception and cognitive reasoning. Efforts in affective perception concentrated on emotion recognition, with models developed for both coarse-grained, utterance-level categories [7, 8] and more nuanced, word-level emotional cues [9, 3]. In parallel, the cognitive reasoning stream aimed to deepen a model’s comprehension of the user’s situation, often by integrating external knowledge sources like commonsense graphs [4, 10]. While these specialized approaches advanced the analytical aspects of empathy, a persistent limitation was the difficulty in generating responses that matched the sophistication of their analytical insights, often due to constrained model parameters [11].
@@ 2.2 Generative Fluency and The Synthesis Challenge
The advent of LLMs such as ChatGPT [12] marked a significant leap in generative fluency for ERG. By leveraging vast pre-training, LLMs excel at producing coherent and contextually appropriate language, overcoming the expressive limitations of earlier models [13, 14]. However, this fluency has not always been accompanied by a corresponding depth of empathetic understanding. Multiple studies indicate that LLMs can falter in fine-grained emotional perception and may not perform the necessary reasoning to uncover the root causes of a user’s feelings. This highlights a trade-off between analytical depth and generative power. We propose TRACE to address this, a multi-agent framework that decomposes the empathetic process into a structured pipeline. By dedicating distinct agents to emotion perception, causal analysis, and strategic planning, our approach is designed to unite the deep, structured understanding of specialized models with the powerful, expressive generation of LLMs.

### v2

@@ Abstract
Empathetic response generation is a crucial task for creating more human-like and supportive conversational agents. However, existing methods face a core trade-off between the analytical depth of specialized models and the generative fluency of Large Language Models (LLMs). To address this, we propose TRACE, Task-decomposed Reasoning for Affective Communication and Empathy, a novel framework that models empathy as a structured cognitive process by decomposing the task into a pipeline for analysis and synthesis. By building a comprehensive understanding before generation, TRACE unites deep analysis with expressive generation. Experimental results show that our framework significantly outperforms strong baselines in both automatic and LLM-based evaluations, confirming that our structured decomposition is a promising paradigm for creating more capable and interpretable empathetic agents. Our code is available at https://github.com/sunbus100/TRACE.
@@ 2.1 Modeling Empathetic Understanding
Early research in Empathetic Response Generation primarily focused on modeling the user’s inner state to build a foundation for empathetic interaction [20, 5]. This pursuit branched into two main research streams, affective perception and cognitive reasoning. Efforts in affective perception concentrated on emotion recognition, with models developed for both coarse-grained, utterance-level categories [13, 18] and more nuanced, word-level emotional cues [8, 12]. In parallel, the cognitive reasoning stream aimed to deepen a model’s comprehension of the user’s situation, often by integrating external knowledge sources like commonsense graphs [22, 3]. While these specialized approaches advanced the analytical aspects of empathy, a persistent limitation was the difficulty in generating responses that matched the sophistication of their analytical insights, often due to constrained model parameters [2].
@@ 2.2 Generative Fluency and The Synthesis Challenge
The advent of LLMs such as ChatGPT [1] marked a significant leap in generative fluency for ERG. By leveraging vast pre-training, LLMs excel at producing coherent and contextually appropriate language, overcoming the expressive limitations of earlier models [29, 27]. However, this fluency has not always been accompanied by a corresponding depth of empathetic understanding. Multiple studies indicate that LLMs can falter in fine-grained emotional perception and may not perform the necessary reasoning to uncover the root causes of a user’s feelings [16]. This highlights a trade-off between analytical depth and generative power. We propose TRACE to address this, a multi-agent framework that decomposes the empathetic process into a structured pipeline. By dedicating distinct agents to emotion perception, causal analysis, and strategic planning, our approach is designed to unite the deep, structured understanding of specialized models with the powerful, expressive generation of LLMs.
@@ 5 CONCLUSION
6 Acknowledgment
This research was supported by the National Natural Science Foundation of China (Grant No. 62502396) and the XJTLU Research Development Fund (RDF-24-02-008).